A ladder is resting against a wall. The top of the ladder touches the wall at the height of 15 feet. Find the length of the ladder if the length of the ladder if the length is 5 feet more than its distance from the wall.

Answer by mananth(16946)

A ladder is resting against a wall. The top of the ladder touches the wall at the height of 15 feet. Find the length of the ladder if the length of the ladder if the length is 5 feet more than its distance from the wall.
The ladder ,ground and the wall form a right triangle
ground distance = leg1
height at which ladder touches
Hypotenuse = ladder length
Pythagoras theorem states that

Hypotenuse ^2= leg1^2+leg2^2
height at which ladder touches wall = 15 ft
length of ladder = x+5
distance from wall = x
15^2 = (x+5)^2-x^2
225 = x^2+10x+25-x^2
225=10x+25
-225
10x-200=0
10x = 200
/10
x= 20 distance from wall
(x+5 )= 25 ft the length of ladder
m.ananth@hotmail.ca